# Photoacoustic Iterative Optimization Algorithm with Shape Prior Regularization

### v2

@@ Abstract
Photoacoustic imaging (PAI) suffers from inherent limitations that can degrade the quality of reconstructed results, such as noise, artifacts and incomplete data acquisition caused by sparse sampling or partial array detection. In this study, we proposed a new optimization method for both two-dimensional (2D) and three-dimensional (3D) PAI reconstruction results, called the regularized iteration method with shape prior. The shape prior is a probability matrix derived from the reconstruction results of multiple sets of random partial array signals in a computational imaging system using any reconstruction algorithm, such as Delay-and-Sum (DAS) and Back-Projection (BP). In the probability matrix, high-probability locations indicate high consistency among multiple reconstruction results at those positions, suggesting a high likelihood of representing the true imaging results. In contrast, low-probability locations indicate higher randomness, leaning more towards noise or artifacts. As a shape prior, this probability matrix guides the iteration and regularization of the entire array signal reconstruction results using the original reconstruction algorithm (the same algorithm for processing random partial array signals). The method takes advantage of the property that the similarity of the object to be imitated is higher than that of noise or artifact in the results reconstructed by multiple sets of random partial array signals of the entire imaging system. The probability matrix is taken as a prerequisite for improving the original reconstruction results, and the optimizer is used to further iterate the imaging results to remove noise and artifacts and improve the imaging fidelity. Especially in the case involving sparse view which brings more artifacts, the effect is remarkable. Simulation and real experiments have both demonstrated the superiority of this method.
@@ 1 Introduction
Photoacoustic imaging (PAI) has emerged as a promising imaging modality that combines the high spatial resolution of ultrasound with the rich optical contrast of biological tissues.[1] This hybrid imaging technique has found wide-ranging applications in biomedical imaging due to its ability to provide detailed structural and functional information, particularly in scenarios such as vascular imaging[2], tumor detection[3], and brain imaging[4]. However, as with many computational imaging systems, PAI suffers from inherent limitations that can degrade the quality of reconstructed results, such as noise, artifacts, and incomplete data acquisition caused by sparse sampling or partial array detection. Traditional image reconstruction methods, while effective to some extent, often struggle to handle such challenges, especially when the imaging system faces conditions like a sparse view or other constraints that exacerbate result artifacts.[5] These approaches may fail to fully exploit the inherent redundancies or structural priors present in the data, leading to suboptimal performance in terms of reconstruction fidelity and artifact suppression.
@@
      To tackle these challenges, researchers have utilized iterative reconstruction (IR) methods to enhance result quality when using sparse view. One of the earliest IR approaches for PAI was introduced by Paltauf et al., who iteratively minimized the discrepancies between observed and simulated projections to improve image reconstruction.[6] Subsequent IR techniques incorporated various physical factors related to imaging. Later, Deán-Ben et al. developed a precise model-based 3D reconstruction algorithm that used a sparse matrix model to compute theoretical pressure[7], resulting in fewer artifacts compared to back-projection methods. Wang et al.[8, 9] demonstrated that iterative penalized least-squares methods, constructed on discrete-to-discrete imaging models with expansion functions over different fixed spatial grids, and using either quadratic smoothing or total variation (TV) norm constraints, could significantly improve the performance of 3D PAI systems for small animals. Similarly, Huang et al.[10] proposed forward and backward operators based on k-space pseudospectral methodologies. While these IR techniques have shown excellent results, their applicability to large-scale 3D PAI is limited due to the substantial increase in computational and memory demands caused by the scaling of the spatial grid and system matrix.[11] Researchers have, therefore, explored novel alternatives. Arridge et al. addressed this by introducing efficient numerical implementations of adjoint operators for PAI reconstruction and applying compressed sensing via Bregman iteration to reduce sensor requirements.[12] However, their method was computationally expensive and restricted to voxel counts on the scale of $10^{6}$. Shang et al.[13] took a different approach by building a forward model using directly measured graphite point sources, but the inefficiencies in optimization and inaccuracies in the model itself limited its practicality. Deep learning-based methods for PAI reconstruction improve computational efficiency and image quality[14, 15, 16] but depend on large, specialized datasets that are hard to obtain and struggle to generalize across different imaging systems or conditions, limiting their clinical applicability.
@@
      In this study, we proposed a new optimization method for photoacoustic imaging reconstruction results. The key idea of this method is that the similarity of the imaged object is higher than the noise and artifacts in results of the reconstruction of multiple random partial array signals in the entire imaging array using any reconstruction algorithm, such as DAS and BP, and using these results a probability matrix is obtained through calculation methods. The different positions of the matrix reflect the probability that the position is the real object in the imaging space. As a shape prior[17], this probability matrix guides the iteration and regularization of the entire array signal reconstruction results using the original reconstruction algorithm (the same reconstruction algorithm for processing random partial array signals). This method is not affected by the array distribution, imaging region, etc., and can be used in both 2D and 3D imaging in photoacoustic computational imaging, and can be widely applied to other computational imaging systems, including ultrasound, ct, etc., with strong generalization and robustness. By leveraging the shape prior based on possibility matrix and iteratively refining reconstruction results, this method achieves superior performance compared to the result before the optimization. Both simulation and experimental results demonstrate the effectiveness of the proposed method in reducing noise and artifacts under sparse-view conditions while preserving critical structural information in PAI reconstructions. For the convenience of calling this algorithm, we call it RI-SPPM, RI means that it is a regularized iteration process, SPPM means that the shape prior is based on a probability matrix.
@@
      This paper is organized as follows. In Section 2, we present the theoretical framework of the proposed method, detailing the construction of the probability matrix, the integration of the shape prior, and the iterative algorithm employed. Section 3 describes the simulation and experimental results, demonstrating the advantages of our method. Finally, Section 4 concludes the paper by discussing the implications of the proposed method and exploring potential future developments for RI-SPPM in computational imaging applications.
@@ 2.1 Construction of Probability Matrix
This section begins by introducing the causes of noise and artifacts in photoacoustic imaging, laying the foundation for the rationale of the proposed method. For the construction of the probability matrix, we randomly select a portion of the signal from all array elements for reconstruction. A probability matrix consistent with the dimensions of the reconstruction results is generated using mathematical operations on the obtained reconstruction results. In this matrix, high-probability regions indicate a higher likelihood of representing the true imaged object, while noise and artifacts, due to their randomness, tend to correspond to regions with lower probabilities. Finally, we describe how this shape prior knowledge is utilized in a specific regularized iterative reconstruction process, including the selection of the loss function and the optimizer.
2.1.1 Theoretical Basis: Causes of Noise and Artifacts
2.1.1.1 Noise
In photoacoustic computational imaging, many kinds of noise may be introduced due to the characteristics of the probe and the working environment, which will affect the image quality. These noises include electronic and thermal noise of the probe itself, frequency distortion due to limited bandwidth, angle-dependent noise due to directivity, mechanical vibration and external acoustic interference, sensitivity or position differences between probe arrays, etc. Among these noises, the electronic thermal noise is generated by electronic components such as resistors in the probe and its signal processing system, and is related to its own temperature. Due to the directivity of the probe, the response to signals from non-optimal angles is weak, resulting in angle-dependent directional noise. The sensitivity error of the array probe is due to the different probe elements in the array may be due to the difference in manufacturing process, resulting in sensitivity and position deviation, resulting in a certain error. Then the noise related to the array elements will appear random with the random selection of the array elements. Therefore, we can take advantage of the fact that the randomness of some noise changes with the array element is higher than that of the imaged object with the array element, so as to remove the noise with high randomness and retain the imaged object with high consistency[18].
2.1.1.2 Artifacts
Recorded photoacoustic (PA) signals need to undergo a mathematical transformation to form final PA result. The transformation is typically achieved using reconstruction algorithms, such as BP, and IR. However, different reconstruction algorithms possess distinct characteristics and may produce images with different artifacts. Common artifacts in PAI, including negativity artifacts, streak artifacts, splitting artifacts, and reflection artifacts, affect the quality of reconstruction results.
Streak Artifacts
As the artifact most affected by the probe distribution, Streak artifacts describe undesired structures in an image that have a streak shape. They are a common type of artifacts in PA imaging and are formed due to non-ideal imaging conditions,such as sparse and limited-angle sampling. Most streak artifacts occur around PA absorbers that have stronger optical absorption than their background. The spatial resolution of PA imaging can be greatly impacted by the streak artifacts, especially when strong PA absorbers are present.Figure 1 shows the results of 2D reconstruction under different array configurations. It can be clearly found that this type of artifact is greatly affected by array configurations.Therefore, we can make good use of the characteristic that the randomness of the streak artifact changing with the array is higher than that of the imaged object changing with the array, so as to remove the streak artifact with high randomness and retain the imaging object with high consistency[18].
Figure 1: Streak artifacts in PAI. (a) A numerical phantom with two disks. (b) and (c) Reconstructed images using BP with 32 and 16 detectors evenly-distributed on a full ring. (d) Reconstructed images by using half ring (16 detectors)
Negativity Artifacts
Compared to streak artifacts, negative artifacts are less influenced by the randomness of matrix distribution but are still impacted by its properties. Negativity artifacts refer to unwanted structures in photoacoustic (PA) results that exhibit negative magnitudes. These artifacts are typically found at the edges of PA sources and are considered undesirable because PA results are intended to represent the optical absorbance characteristics of biological tissues, which are inherently positive. Research has indicated that two primary factors contributing to these artifacts are limited detector bandwidth and restricted view angle. When the detector’s bandwidth is insufficient, the back-projected (BP) signals become distorted due to the loss of essential frequency components, leading to the emergence of negativity artifacts. Similarly, with a limited view angle, certain PA signals propagating in three-dimensional space may not be captured, resulting in partial data loss and artifact formation. Additionally, a non-ideal acoustic medium can further exacerbate negativity artifacts, as acoustic attenuation, scattering, and refraction may distort the signals. These artifacts are not exclusive to BP algorithms but are also prevalent in other reconstruction approaches, such as time-reversal (TR) and frequency-domain methods.As shown in Figure 2, the array distribution has a certain influence on the formation position of negative artifacts.Therefore, we can make good use of the characteristic that the randomness of the negativity artifact changing with the array is higher than that of the imaged object changing with the array, so as to remove the artifact with high randomness and retain the imaging object with high consistency[18].
Figure 2: Negativity artifacts schematic diagram(blue is the negative artifact)
Other Artifacts
For other artifacts, there is no detailed research report or scientific data to show the influence of array properties such as array distribution on artifacts, such as Splitting artifacts, Reflection artifacts, etc., but in general, compared to the real objects to be imitated, Artifacts are intuitively more strongly influenced by element randomness
@@ 2.1.2 Reconstruction of Random Selected Element Signals
For a computational imaging system with array number $N$ and sampling number $M$, the obtained signal domain data dimension is $N*M$. We randomly select $x$ element signals for reconstruction ($x<N$), here we have A selections(A = $c_{N}^{x}$), and we only need A random k results ($k<A$), the $k$ results are $R_{1},R_{2},...R_{k}$. (The choices of $x$ and $k$ are described in detail in subsequent paragraphs.)
2.1.3 Acquisition of Probability Matrix
In three dimensions, the probability matrix is defined as $P(x,y,z)$.
$P(x,y,z)=\int_{v_{\min}}^{v_{\max}}D(x,y,z)\cdot\frac{v-v_{\min}}{v_{\max}-v_{\min}}\cdot\|D(x,y,z)-v>0\|_{0}\mathrm{d}v$ (1)
where, $v_{min}=min(D(x,y,z))$, $v_{max}=max(D(x,y,z))$
$D(x,y,z)=\frac{(\frac{1}{k}(R_{1}+R_{2}+...+R_{k}))^{2}}{\frac{1}{k}(R_{1}^{2}+R_{2}^{2}+...R_{k}^{2})}=\frac{(R_{1}+R_{2}+...+R_{k})^{2}}{{k}\times(R_{1}^{2}+R_{2}^{2}+...R_{k}^{2})}$ (2)
In order to have a more intuitive physical meaning for the subsequent regular iteration process, we normalize $P$ $0-1$. The result after normalization is $P_{N}$.
${P_{N}=\frac{P-min(P)}{max(P)-min(P)}}$ (3)
The function $P(x,y,z)$ is introduced as a probabilistic representation of spatial data distributions in three-dimensional space. The function integrates the local property $D(x,y,z)$, which quantifies the concentration or uniformity of measured attributes $R_{1},R_{2},...,R_{k}$ at a given point $(x,y,z)$, defined as the ratio of the squared arithmetic mean to the mean of squared values. This makes $D(x,y,z)$ a measure of localized data consistency, where higher values indicate greater uniformity. To account for variations across the defined region, $P(x,y,z)$ is computed by integrating over the range $[v_{m}in,v_{m}ax]$, where these bounds represent the global extremal values of $D(x,y,z)$. The integration involves a weighting factor $\frac{v-v_{\min}}{v_{\max}-v_{\min}}\cdot\|D(x,y,z)-v\|_{0}$, which characterizes the probabilistic "importance" of this point within the entire three-dimensional region through weighting and integration range.
@@ 2.2 Regularized Iteration Optimization
In this section, we will use the probability matrix $P$ obtained in the previous section as a shape prior and design a suitable loss function for regular iterative optimization of the original reconstruction results of the entire array signal.
@@ 2.2.1 Shape Prior and Loss Function
2.2.1.1 Shape Prior
The feasibility of using $P_{N}$ as a shape prior lies in the fact that it directly provides the probability distribution information of each voxel or pixel region, which can be used to indicate the model’s tendency to strengthen certain regions (high probability regions) in the reconstruction results and its inhibition demand for other regions (low probability regions). This prior information can effectively guide the optimization process, making the result more in line with the expected structural characteristics, while reducing the impact of noise or artifacts. The rational constraint of shape information is realized mathematically, and the optimization stability of shape consistency is improved. It has clear feasibility and physical significance.
2.2.1.2 Loss Function
The result of reconstructing all array signals using the algorithm to be optimized is $R_{N}$. The result of each iteration optimization is $R_{op}$. To enhance the generalizability of the algorithm and reduce the need for extensive parameter adjustments during application, $R_{N}$ and $R_{op}$ are normalized to the range of $[-1,1]$ during the data preprocessing stage. The loss function mainly consists of two parts: data consistency item[19] and regularized term[20].
Data Consistency Item
Loss of data consistency item is represented by $Loss_{dc}$.
${Loss_{dc}}=(R_{N}-R_{op})^{2}$ (4)
@@ 2.2.1.2 Loss Function
${gradient_{dc}=\frac{\partial{Loss}_{dc}}{\partial R_{op}}=2\times(R_{N}-R_{op})}$ (5)
The role of the data consistency item is to ensure that the optimized results do not deviate too much from the original reconstruction results, thereby maintaining physical consistency and data authenticity with the input data. In the photoacoustic computational reconstruction problem, the result $R_{N}$ reconstructed by the initial method to be optimized is obtained by a certain inversion method. Although it may contain noise or error, it is still the basic data closest to the real physical field distribution. Due to the regularization introduced in the optimization process, the $R_{o}p$ of the optimization result may deviate from the original data, and even lose the important features of the original signal. Therefore, the data consistency item can impose constraints in the optimization process, forcing the optimization results to be consistent with the original reconstructed data as much as possible, so that the optimized volume data can gradually de-noise and enhance the structure while still reflecting the main features and information of the input data. This is a balancing mechanism that avoids distortion caused by excessive regularization and strengthens the physical confidence and data integrity of the optimized results.
Regularized Term
Loss of regularized term is represented by $gradient_{rg}$.
${Loss_{rg}}=((1-P_{N})\times R_{op})^{2}$ (6)
@@
${gradient_{rg}=\frac{\partial{Loss}_{rg}}{\partial R_{op}}=2\times(1-P_{N})\times R_{op}}$ (7)
The significance of regularization term is to use the prior knowledge of the problem to constrain the optimization process. As we mentioned in the previous section, in the actual reconstruction, the results obtained by using the traditional reconstruction algorithm are accompanied by noise and artifacts, which makes the reconstruction results rough and even difficult to reflect the real shape of the target. By combining the prior information provided by the prior probability distribution $P_{N}$, the regularization term can actively suppress the data in the low probability region, thereby reducing the interference of artifacts or noise to the solution in the reconstruction process. The design idea of the regularization term is to guide the optimal solution to cluster in the high probability region (these regions are more in line with the prior physical knowledge), that is, the penalty of the high probability region is low, while the value of the low probability region is suppressed, that is, the penalty of the low probability region is high. This constraint can make the optimization more robust, that is, reduce overly radical updates in areas with greater uncertainty, and ultimately promote the optimized results to have both prior rationality and data authenticity, thus generating reconstruction results with higher reliability.

### v4

# Iterative Optimization Algorithm with Shape Prior to Reduce Noise and Artifacts for Photoacoustic Imaging

@@ Abstract
Although exact reconstruction algorithms exit for Photoacoustic imaging (PAI), in reality its image quality is inevitably influences by system noise and artifacts. In this study, we propose an iterative optimization method for PAI reconstruction, called the regularized iteration method with shape prior (RISP). The shape prior is a probability matrix derived from multiple reconstruction results via randomly partial array elements using any reconstruction algorithms, such as widely used Delay-and-Sum (DAS) and Filtered Back-Projection (FBP). In the probability matrix, high-probability values indicate high coherency among multiple reconstruction results at those positions, suggesting a high likelihood of representing the true imaging results. In contrast, low-probability values indicate higher randomness, leaning more towards noise or artifacts. As a shape prior, this probability matrix, together with the orginal PAI result using all array elements, guides the iteration and regularization of the PAI results. The probability matrix is taken as a prerequisite for improving the original reconstruction results, and the optimizer is used to further iterate the imaging results to remove noise and artifacts and improve the imaging fidelity. The simulation and real PAI animal and human results demonstrated our method can substantially reduce both noise and artefacts without the need of hardware upgrade.
@@ 1 Introduction
PAI has emerged as a promising imaging modality that combines the high spatial resolution of ultrasound with the rich optical contrast of biological tissues[1]. This hybrid imaging technique has found wide-ranging applications in biomedical imaging due to its ability to provide detailed structural and functional information, particularly in scenarios such as vascular imaging[2], tumor detection[3], and brain imaging[4]. However, in reality, the image quality of PAI generally suffers from inherent limitations that can degrade the quality of reconstructed results, such as noise, artifacts, and incomplete data acquisition caused by sparse sampling or partial array detection. Traditional image reconstruction methods, while effective to some extent, often struggle to handle such challenges, especially when the imaging system faces conditions like a sparse view or other constraints that exacerbate result artifacts.[5] These approaches may fail to fully exploit the inherent redundancies or structural priors present in the data, leading to suboptimal performance in terms of reconstruction fidelity and artifact suppression.
@@
      To tackle these challenges, researchers have utilized iterative reconstruction (IR) methods to enhance result quality when using sparse view. One of the earliest IR approaches for PAI was introduced by Paltauf et al., who iteratively minimized the discrepancies between observed and simulated projections to improve image reconstruction.[6] Subsequent IR techniques incorporated various physical factors related to imaging. Later, Deán-Ben et al. developed a precise model-based 3D reconstruction algorithm that used a sparse matrix model to compute theoretical pressure[7], resulting in fewer artifacts compared to back-projection methods. Wang et al.[8, 9] demonstrated that iterative penalized least-squares methods, constructed on discrete-to-discrete imaging models with expansion functions over different fixed spatial grids, and using either quadratic smoothing or total variation (TV) norm constraints, could significantly improve the performance of 3D PAI systems for small animals. Similarly, Huang et al.[10] proposed forward and backward operators based on k-space pseudospectral methodologies. While these IR techniques have shown excellent results, While these IR techniques have shown excellent results, the computing burden in both time and memory consumption is still a serious challenge for their implementations, especially for large-scale 3D PAI.[11] Researchers have, therefore, explored novel alternatives. Arridge et al. addressed this by introducing efficient numerical implementations of adjoint operators for PAI reconstruction and applying compressed sensing via Bregman iteration to reduce sensor requirements.[12] However, their method was computationally expensive and restricted to voxel counts on the scale of $10^{6}$. Shang et al.[13] took a different approach by building a forward model using directly measured graphite point sources, but the inefficiencies in optimization and inaccuracies in the model itself limited its practicality. Deep learning-based methods for PAI reconstruction improve computational efficiency and image quality[14, 15, 16] but depend on large, specialized datasets that are hard to obtain and struggle to generalize across different imaging systems or conditions, limiting their clinical applicability.
@@
      Unlike the traditional IR algorithms that demands repeating solving forward model for every iteration circle, here, we proposed a new optimization method for the results of photoacoustic imaging reconstruction that does not need solving the forward model. The key idea of this method is that the similarity of the structure representing the true imaged object is much higher than the random noise and variation artifacts highly related to the arrangement of sensor positions and numbers. Therefore, we construct a probability matrix based on a large number of image reconstruction results from sub-sets of randomly picked original array elements using any reconstruction algorithm, such as DAS and BP. And the probability matrix is obtained through calculation methods. The different positions of the matrix reflect the probability that the position is the real object in the imaging space. As a shape prior[17], this probability matrix guides the iteration and regularization of the originally reconstructed results using total array elements. This method is not affected by specific reconstruction algorithms, such as DAS and UBP, and it can be applied for any array configurations. By leveraging the shape prior based on possibility matrix and iteratively refining reconstruction results, this method substantially improved image quality while keeping quantitative accuracy. Both simulation and experimental results demonstrate the effectiveness of the proposed method in reducing noise and artifacts under sparse-view conditions while preserving critical structural information in PAI reconstructions.
@@
      This paper is organized as follows. In Section 2, we present the theoretical framework of the proposed method, detailing the construction of the probability matrix, the integration of the shape prior, and the iterative algorithm employed. Section 3 describes the simulation and experimental results, demonstrating the advantages of our method. Finally, Section 4 concludes the paper by discussing the implications of the proposed method and exploring potential future developments for RISP in computational imaging applications.
@@ 2.1.1 Reconstruction of Random Selected Element Signals
For a PAI array system with the element number $N$ and temporal sampling number $M$, the data dimension of recorded signal is $N\times M$. Besides using all $N$ elements to do image reconstruction, we randomly select $x$ ($x<N$) sub-elements for reconstruction. The total number of selection is A (A = $c_{N}^{x}$), and we only randomly pick k results ($k<A$), naming $R_{1},R_{2},...,R_{k}$. (The choices of $x$ and $k$ are described in detail in subsequent paragraphs.)
2.1.2 Construction of Probability Matrix
In 3D PAI, the probability matrix is defined as $P(x,y,z)$.
$P(x,y,z)=D(x,y,z)\odot\int_{v_{\min}}^{v_{\max}}\Phi(v|\mu,\sigma)\odot\mathrm{u}(D(x,y,z)-v)\mathrm{d}v$ (1)
where, $\mathrm{u}(x)$ is the unit step function, and $D(x,y,z)$ is calculated by
$D(x,y,z)=\frac{(\frac{1}{k}(R_{1}+R_{2}+\cdots+R_{k}))^{2}}{\frac{1}{k}(R_{1}^{2}+R_{2}^{2}+\cdots+R_{k}^{2})}=\frac{(R_{1}+R_{2}+\cdots+R_{k})^{2}}{k(R_{1}^{2}+R_{2}^{2}+\cdots+R_{k}^{2})}$ (2)
$v_{min}=\mathrm{min}(D(x,y,z)),\ v_{max}=\mathrm{max}(D(x,y,z)),\ \mu=(v_{max}+v_{min})/2,\ \sigma=(v_{max}-v_{min})/6$, and $\Phi(v|\mu,\sigma)$ is the cumulative distribution function of the normal distribution.
$\Phi(v|\mu,\sigma)=\frac{1}{\sigma\sqrt{2\pi}}\int_{-\infty}^{v}e^{\frac{-(t-\mu)^{2}}{2\sigma^{2}}}\mathrm{d}t$ (3)
In order to have a more intuitive physical meaning for the subsequent regular iteration process, we use normalized $P_{norm}$ $(0<P_{norm}<1)$.
${P_{norm}=\frac{P-\mathrm{min}(P)}{\mathrm{max}(P)-\mathrm{min}(P)}}$ (4)
The Matrix $P_{norm}(x,y,z)$ is a probabilistic representation of spatial data distributions. The function integrates the local property $D(x,y,z)$, which quantifies the concentration or uniformity of measured attributes $R_{1},R_{2},...,R_{k}$ at a given spatial point $(x,y,z)$, defined as the ratio of the squared arithmetic mean to the mean of squared values. This makes $D(x,y,z)$ a measure of localized data consistency, where higher values indicate greater consistency. The introduction of the Gaussian cumulative distribution-based weighting factor $\Phi(v|\mu,\sigma)$ combines local signal intensities with global distribution characteristics, ensuring that the weight assignment is both scientifically robust and consistent. The Gaussian distribution, determined by its mean $\mu$ and standard deviation $\sigma$, dynamically reflects the global distribution pattern of signals. Strong signal regions above the mean receive higher weights, while weaker signals are appropriately suppressed but not entirely discarded. This smoothing property avoids the adverse effects of outliers or noisy signals on the results while presenting a clear and hierarchical structure of signal importance across the global range. Consequently, this weighting approach enhances signal saliency, suppresses background noise, and selectively preserves significant regions, achieving more robust and efficient reconstruction. This integral effectively blends the local signal characteristics with global weighting, enabling the assignment of weights that emphasize real signals while preserving the overall smooth and consistent distribution across the signal space.
@@ 2.2 Regularized Iteration Optimization
In this section, we will use the probability matrix $P_{norm}$ obtained in the previous section as a shape prior and design a suitable loss function for regular iterative optimization of the original reconstruction results of the entire array signals.
@@ 2.2.1 Shape Prior and Loss Function
Shape Prior
The feasibility of using $P_{norm}$ as a shape prior lies in the fact that it directly provides the probability distribution information of each voxel or pixel region, which can be used to indicate the model’s tendency to strengthen certain regions (high probability regions) in the reconstruction results and its inhibition demand for other regions (low probability regions). This prior information can effectively guide the optimization process, making the result more in line with the expected structural characteristics, while reducing the impact of noise or artifacts. The rational constraint of shape information is realized mathematically, and the optimization stability of shape consistency is improved. It has clear feasibility and physical significance.
Loss Function
The result of reconstructing all array signals using the algorithm to be optimized is $R_{N}$. The result of each iteration optimization is $R_{op}$. To enhance the generalizability of the algorithm and reduce the need for extensive parameter adjustments during application, $R_{N}$ and $R_{op}$ are normalized to the range of $[0,1]$ after taking the absolute value operation. The loss function consists of two terms: data consistency loss term[18] and regularization loss term[19].
Data Consistency Loss Term
The data consistency loss term is represented by $Loss_{dc}$.
${Loss_{dc}}=||R_{N}-R_{op}||^{2}_{F}$ (5)
@@
${gradient_{dc}=\frac{\partial{Loss}_{dc}}{\partial R_{op}}=2\odot(R_{N}-R_{op})}$ (6)
The role of the data consistency loss term is to ensure that the optimized results do not deviate too much from the original reconstruction results, thereby maintaining physical consistency and quantitative accuracy. Although $R_{N}$ contains noise or artifacts, it is still close to the true image. Due to the regularization introduced in the optimization process, the $R_{o}p$ of the optimization result may deviate from the original data, and even lose the important features of original signals. Therefore, the data consistency loss term imposes constraints in the optimization process, forcing the optimization results to be consistent with the original reconstructed data as much as possible, so that the optimized volume data can gradually de-noise and enhance the structure while still reflecting the main features and information of the input data. This is a balancing mechanism that avoids distortion caused by excessive regularization and strengthens the physical confidence and data integrity of the optimized results, which is essential for quantitative accuracy.
Regularization Loss Term
The regularization loss term is represented by $Loss_{rg}$.
${Loss_{rg}}=||(1-P_{N})\odot R_{op}||^{2}_{F}$ (7)
@@
${gradient_{rg}=\frac{\partial{Loss}_{rg}}{\partial R_{op}}=2\times(1-P_{N})\odot R_{op}}$ (8)
The significance of regularization loss term is to use the shape prior knowledge to constrain the optimization process. As we mentioned in the previous section, in the actual reconstruction, the results obtained by using the traditional reconstruction algorithm are accompanied by noise and artifacts, which makes the reconstruction results bad and even difficult to reflect the real shape of the target. By combining the prior information provided by the prior probability distribution $P_{N}$, the regularization term can actively suppress the data in the low probability region, thereby reducing the interference both of artifacts and noise to the solution in the reconstruction process. The design idea of the regularization term is to guide the optimal solution to cluster in the high probability region (these regions are more in line with the prior physical knowledge), that is, the penalty of the high probability region is low, while the value of the low probability region is suppressed, that is, the penalty of the low probability region is high. This constraint can make the optimization more robust, that is, reduce overly radical updates in areas with greater uncertainty, and ultimately promote the optimized results to have both prior rationality and data authenticity, thus generating reconstruction results with higher reliability.